Robin's company invested its profits of $52,700, part at 13.5% and part at 15%. the total annual yield on this investment was $7432.50. How much was invested at each rate?
Answer by checkley77(12844)

.15X+.135(52,700-X)=7,432.50
.15X+7,114.5-.135X=7,432.50
.015X=7,432.50-7,114.50
.015X=318
X=318/.015
X=21,200 ANSWER FOR THE AMOUNT INVESTED @ 15%.
52,700-21,200=31,500 ANSWER FOR THE 13.5$ INVESTMENT.
PROOF:
.15*21,200+.135*31,500=7,432.50
3,180+4,252.5=7,432.5
7,432.5=7,432.5